OpenStudy (anonymous):

The table below shows two equations: Equation 1 |3x – 1| + 7 = 2 Equation 2 |2x + 1| + 4 = 3 Which statement is true about the solution to the two equations? Equation 1 and equation 2 have no solutions. Equation 1 has no solution and equation 2 has solutions x = 0, 1. The solutions to equation 1 are x = -1.3, 2 and equation 2 has no solution. The solutions to equation 1 are x = -1.3, 2 and equation 2 has solutions x = 0, 1.

OpenStudy (anonymous):

hlep

OpenStudy (math&ing001):

Simplify them like this : |3x - 1| = -5 |2x + 1| = -1 we know that |.| is always positive, what can you deduce ?

OpenStudy (anonymous):

is b

OpenStudy (math&ing001):

What's the reason ?

OpenStudy (anonymous):

dont know please help

OpenStudy (math&ing001):

|3x - 1| and |2x + 1| are both positive and -1 and -5 both negative. A positive number can not equal a negative number. Do you get what I mean ?

OpenStudy (anonymous):

what is that anwer please or help

OpenStudy (math&ing001):

We don't just give answers here, we guide people to the answer. You have to make an effort on your part, otherwise I can't help you.

OpenStudy (texaschic101):

| 3x - 1 | + 7 = 2 --- subtract 7 from both sides | 3x - 1 | = 2 - 7 | 3x - 1 | = - 5 (3x - 1) = -5 3x = -5 + 1 3x = -4 x = -4/3 which equals -1.3 -(3x - 1) = -5 -3x + 1 = -5 -3x = -5 - 1 -3x = -6 x = 2 solution to 1 is : 2 and -1.3 | 2x + 1| + 4 = 3 | 2x + 1 | = 3 - 4 | 2x + 1 | = -1 (2x + 1) = -1 2x = -1-1 2x = -2 x = -1 -(2x + 1) = -1 -2x - 1 = -1 -2x = -1 + 1 -2x = 0 x = 0 solution to 2 is : -1 and 0 answer to question is D

OpenStudy (math&ing001):

@texaschic101 That's unfortunately not the correct answer. |2x + 1| and |3x - 1| are both positive. And since -5 and -1 are negative, both equations don't have a solution.